1.  y'' + 4y' + 7y = 0         ans, y= e-2x(c1cos"sqr3" x+c2sin"sqr3" x)

1
Expert's answer
2021-12-10T08:50:50-0500

1.

Corresponding (auxiliary) equation


"r^2+4y+7=0"

"D=(4)^2-4(1)(7)=-12"

"r=\dfrac{-4\pm \sqrt{12}}{2}=-2\pm i\sqrt{3}"

The general solution of the given differential equation is


"y=c_1e^{-2x}\cos(\sqrt{3}x)+c_2e^{-2x}\sin(\sqrt{3}x)"